# Classification of coadjoint orbits for the maximal unipotent subgroup in the simple group of type $F_{4}$

### v3

@@ Abstract
Let NN be the maximal unipotent subgroup in the simple algebraic group of type $\Phi$. It naturally acts on the space $\mathfrak{n}^{*}$ dual to the Lie algebra $\mathfrak{n}$ of NN, and this action is called coadjoint. Such orbits play the key role in the A.A. Kirillov’s orbit method. In this work, we classify the orbits of this action in the case of $\Phi=F_{4}$ in terms of supports of canonical forms. This means that we will present a set $S$ of linear forms from $\mathfrak{n}^{*}$ such that for any coadjoint orbit there exists a unique form from $S$ belonging to that orbit. The set of canonical forms will be explicitly described in terms of supports. The support of a form $\lambda\in\mathfrak{n}^{*}$ is the set of positive roots $\alpha\in\Phi^{+}$ such that $\lambda(e_{\alpha})\neq 0$.
@@ 1. Introduction and the main result
Let $\mathfrak{g}$ be a simple finite-dimensional Lie algebra over a finite field $\mathbb{F}_{q}$ of sufficiently large characteristic $p$, $\mathfrak{b}$ be a Borel subalgebra in $\mathfrak{g}$, $\Phi$ be the root system of $\mathfrak{g}$, $\Phi^{+}$ be the set of positive roots corresponding to $\mathfrak{b}$, $\mathfrak{n}$ be the nilradical of $\mathfrak{b}$, $N=\exp(\mathfrak{n})$ be the corresponding unipotent algebraic group, and $\mathfrak{n}^{*}$ be the dual space to $\mathfrak{n}$. The group NN is also a maximal unipotent subgroup in the simple algebraic group $G$ of type $\Phi$, or equivalently a Sylow $p$-subgroup in $G$. The group NN acts on $\mathfrak{n}$ by the adjoint action; the dual action of the group NN on the space $\mathfrak{n}^{*}$ is called the coadjoint action; we will denote the result of this action by $g.\lambda$ for $g\in N$, $\lambda\in\mathfrak{n}^{*}$. According to the orbit method introduced by A.A. Kirillov in 1962, coadjoint orbits play the key role in the representation theory of the group NN. Almost all coadjoint orbits that have been studied well so far are so-called orbits associated with rook placements. They are discussed in detail in the works [An95], [AN06], [Ko12], [Ko13], [Pa08], [IPa09], [Ig09], [Ig11], [Ig12], [IS21] and [IS23].
Explicitly, the coadjoint action looks as follows:
@@
Let $\{e_{\alpha},\alpha\in\Phi^{+}\}$ be a basis in $\mathfrak{n}$ consisting of root vectors, and $\{e_{\alpha}^{*},\alpha\in\Phi^{+}\}$ be the dual basis in $\mathfrak{n}^{*}$. The support of a linear form $\lambda\in\mathfrak{n}^{*}$ is defined as the subset of the set of positive roots of the form
@@
Our goal is to describe the coadjoint orbits in the case of $\Phi=F_{4}$. The description will be given in terms of supports of canonical forms. A list of supports will be provided with the following properties.
  • For any two linear forms $\lambda_{1}\neq\lambda_{2}$ with supports from the given list (possibly, having the same support), the orbits of these forms $N.\lambda_{1}$ and $N.\lambda_{2}$ do not coincide.
  • For any linear form $\lambda\in\mathfrak{n}$, there exists a form $\lambda^{\prime}$ with a support from the given list such that $N.\lambda=N.\lambda^{\prime}$.
In other words, there exists a unique linear form with a support from the given list on each coadjoint orbit.
The set of simple roots in $F_{4}$ can be identified with the following subset of $\mathbb{R}^{4}$:
@@
Here $\{\varepsilon_{i}\}_{i=1}^{4}$ is the standard basis of $\mathbb{R}^{4}$ (with the standard scalar product).
Recall that there exists the natural order on $\Phi^{+}$ defined as follows: $\alpha>\beta$ if $\alpha-\beta$ is a sum of positive roots. Let us fix a linear lexicographic order $\succ$ on the positive roots of $F_{4}$. We will say that $\alpha=a_{1}\alpha_{1}+a_{2}\alpha_{2}+a_{3}\alpha_{3}+a_{4}\alpha_{4}\succ\beta=b_{1}\alpha_{1}+b_{2}\alpha_{2}+b_{3}\alpha_{3}+b_{4}\alpha_{4}$ if $a_{1}>b_{1}$ or $a_{1}=b_{1},a_{2}>b_{2}$ or $a_{1}=b_{1},a_{2}=b_{2},a_{3}>b_{3}$ or $a_{1}=b_{1},a_{2}=b_{2},a_{3}=b_{3},a_{4}>b_{4}$. Thus, the roots will be ordered as follows:
@@
This theorem generalizes the result of the work [IS23]. This article discusses the so-called orbits associated with rook placements. By definition, a subset $D$ of $\Phi^{+}$ is called a non-singular rook placement if $\alpha-\beta\notin\Phi^{+}$ for all distinct $\alpha,\beta\in D$. From this it follows that $(\alpha,\beta)\leq 0$ for all distinct $\alpha,\beta\in D$. Given a rook placement $D$ and a map $\xi\colon D\to\mathbb{C}^{\times}$, we put
@@
We say that the coadjoint orbit $\Omega_{D,\xi}$ of the linear form $f_{D,\xi}$ is associated with the rook placement $D$ and the map $\xi$. In essence, $f_{D,\xi}$ is just a linear form with support $D$.
The work [IS23] proves that if $\Phi=F_{4}$, $D$ is an orthogonal non-singular rook placement and $\xi_{1}$, $\xi_{2}$ are distinct maps from $D$ to $\mathbb{C}^{\times}$ then $\Omega_{D,\xi_{1}}\neq\Omega_{D,\xi_{2}}$. But in fact if $D$ is a non-singular rook placement then $\lambda=f_{D,\xi}\in S$. Because if $\gamma\in D$ then $\gamma+\beta\notin D$ for all $\beta\in\Phi^{+}$ and $\lambda([e_{\gamma},e_{\beta}])=0$ for all $\beta\in\Phi^{+}$. Hence, in the matrix $A_{\lambda,\gamma}$ the line with the number $\gamma$ consists of zeros, and condition (1) is true for $\gamma\in D$. It follows from this that for distinct linear forms with support $D$ the corresponding coadjoint orbits do not coincide.
The result is similar to Theorem Classification of coadjoint orbits for the maximal unipotent subgroup in the simple group of type $F_{4}$ is true for the root system $\Phi=G_{2}$ and field $\mathbb{C}$. In the same work [IS23] the equations of the orbits of basic subvarieties $\mathcal{O}_{D,\xi}$ were written out. By definition, the basic subvariety $\mathcal{O}_{D,\xi}$ corresponding to a rook placement $D$ and a map $\xi\colon D\to\mathbb{C}^{\times}$ is
@@
where $\xi_{\alpha}$ is the restriction of $\xi$ to $\{\alpha\}$. The work also proved the Andrè’s stratification for $G_{2}$:
@@
For almost every rook placement $D$ it is true that $\mathcal{O}_{D,\xi}=\Omega_{D,\xi}$. This is only wrong for $D=\{\alpha+\beta,3\alpha+\beta\}$. In this case $\mathcal{O}_{D,\xi}$ is union of orbits $\Omega_{D,\xi}$ and $\Omega_{\{\beta,\alpha+\beta,3\alpha+\beta\},\xi^{\prime}}$ where $\xi^{\prime}\colon\{\beta,\alpha+\beta,3\alpha+\beta\}\to\mathbb{C}^{\times}$ and $\xi^{\prime}|_{D}=\xi$. Now, let us write out the equations of the orbits for $D=\{\alpha+\beta,3\alpha+\beta\}$. For a linear form $\lambda=\sum_{\gamma\in\Phi^{+}}\lambda_{\gamma}e_{\gamma}^{*}\in\Omega_{D,\xi}$ we have the following system of equations.
@@
The orbits of these forms correspond to all orbits whose equations are written out in the article [IS23] and here.
@@
Acknowledgments. I thank M.V. Ignatev and A.N. Panov for helpful comments and support during writing the paper.
@@ 2. Description of the set of canonical forms
In this section, we will obtain an explicit description of the set $S$ to prove Theorem Classification of coadjoint orbits for the maximal unipotent subgroup in the simple group of type $F_{4}$ and its use. At first glance, it seems that it does not have a good description. In fact, this is not the case. The set $S$ is very well described in terms of supports. In the Appendix, a list of 880 supports will be presented.
@@
whose coordinates satisfy the equation
@@
Here, we mean the coordinates of linear forms in the dual basis, $\lambda=\sum\limits_{\alpha\in\Phi^{+}}\lambda_{\alpha}e_{\alpha}^{*}$.
@@
In the sequel, we will use the following table. It is a $24\times 24$ table, its rows and columns are numbered by positive roots. We write $\beta-\alpha$ in the cell $(\alpha,\beta)$ if $\beta-\alpha$ is the sum of positive roots, otherwise the cell is empty. The main diagonal contains zeros, it is easy to see that all cells below will be empty. Below we present the upper left corner of this table indexed by the positive roots whose decomposition does not contain $\alpha_{1}$. If at position $(\alpha,\beta)$ the difference $\beta-\alpha$ is also a positive root, this cell is highlighted in orange. Almost all examples will be considered specifically on this part of the table. Four columns are highlighted in blue because they will be encountered most often in the examples.
@@
To describe the set $S$, we wrote a program, which is included in the Appendix. The program output supports that satisfy a sufficient condition for being included in the list, or supports that satisfy a necessary condition, these will be described below.
First, let us reformulate condition (1). The algebra $\mathfrak{n}$ acts on $\mathfrak{n}^{*}$ as follows: let $x,y\in\mathfrak{n},\lambda\in\mathfrak{n}^{*}$, then $(x.\lambda)(y):=-\lambda([x,y])$. This action is called tangent.
Let us expand $x\in\mathfrak{n}$ in the basis:
@@
We can consider these coordinates as linear combinations of variables $x_{\alpha},\alpha\in\Phi^{+}$, with constant coefficients $\lambda([e_{\alpha},e_{\beta}])$. By definition, $A_{\lambda,\gamma}$ is the matrix of linear combinations $\{(x.\lambda)(e_{\alpha})\}_{\alpha\succeq\gamma}$, and $B_{\lambda,\gamma}$ is the matrix of linear combinations $\{(x.\lambda)(e_{\alpha})\}_{\alpha\succ\gamma}$. Clearly, $\mathrm{rk}\,A_{\lambda,\gamma}=\mathrm{rk}\,B_{\lambda,\gamma}$ if and only if the row numbered $\gamma$ of the matrix $A_{\lambda,\gamma}$ is a linear combination of all its other rows (i.e., the rows of the matrix $B_{\lambda,\gamma}$). In other words, $(x.\lambda)(e_{\gamma})$ can be expressed as a linear combination of $\{(x.\lambda)(e_{\alpha})\}_{\alpha\succ\gamma}$.
Example 2.2. Let $D=\textup{supp}(\lambda)=\{\alpha_{4},\alpha_{3}+\alpha_{4},\alpha_{2},\alpha_{2}+2\alpha_{3}\}$. We will show that such $\lambda$ lies in $S$. To visualize how the orbit of the tangent action looks like, let us refer to the table. The highlighted blue columns correspond to the roots in $D$. The linear combination $(x.\lambda)(e_{\beta})$ contains only those $x_{\alpha}$ for which $\lambda([e_{\alpha},e_{\beta}])\neq 0$, this is equivalent to $\alpha+\beta\in D$. It can be seen from the construction of the table that $(x.\lambda)(e_{\beta})$ contains only $x_{\alpha}$ such that the orange root $\alpha$ is in the row numbered $\beta$ in the highlighted column. For example, the combination $(x.\lambda)(e_{\alpha_{3}})$ contains only $x_{\alpha_{4}}$ and $x_{\alpha_{2}+\alpha_{3}}$.
@@
It is clear that we did not need to know the coefficients for $x_{\alpha_{3}}$ in either $(x.\lambda)(e_{\alpha_{2}+\alpha_{3}})$ or $(x.\lambda)(e_{\alpha_{4}})$. It is just because $x_{\alpha_{3}}$ as a linear combination was expressed through $(x.\lambda)(e_{\alpha_{2}+\alpha_{3}})$, and everywhere above (i.e., in $(x.\lambda)(e_{\alpha_{2}})$, $(x.\lambda)(e_{\alpha_{3}+\alpha_{4}})$, $(x.\lambda)(e_{\alpha_{3}})$, $(x.\lambda)(e_{\alpha_{4}})$) $x_{\alpha_{3}}$ can be expressed as $(x.\lambda)(e_{\alpha_{2}+\alpha_{3}})$ multiplied by some coefficient. The Sufficient condition below is based on this idea.
@@
Consider the tangent action $\mathfrak{n}\curvearrowright\mathfrak{n}^{*}$. Let $\gamma\in\Phi^{+}$, $\lambda\in\mathfrak{n}^{*}$, $D=\textup{supp}(\lambda)$. Starting from the support $D$ and the root $\gamma$, we will construct the set $\zeta_{D,\gamma}$ consisting of positive roots. We will add roots to $\zeta_{D,\gamma}$ one by one. We will add a root $\delta$ to $\zeta_{D,\gamma}$ if there exists a root $\beta\succ\gamma$ such that $\lambda([e_{\delta},e_{\beta}])\neq 0$ and if $\lambda([e_{\alpha},e_{\beta}])\neq 0$ for some other $\alpha\in\Phi^{+}$ then $\alpha\in\zeta_{D,\gamma}$. Using this procedure, we will add to $\zeta_{D,\gamma}$ the maximum possible number of roots. Note that $\zeta_{D,\gamma}$ does not depend on the values of the coordinates of the form $\lambda$ in the dual basis, but only on its support $D$. The set $\zeta_{D,\gamma}$ also does not depend on the order in which to add the roots. Indeed, let us define the sets $R_{n}$ for $n\in\mathbb{N}$. The root $\delta$ lies in $R_{n}$ if there exists a root $\beta\succ\gamma$ such that $\lambda([e_{\delta},e_{\beta}])\neq 0$ and if $\lambda([e_{\alpha},e_{\beta}])\neq 0$ for some other $\alpha\in\Phi^{+}$ then $\alpha\in R_{i}$ for $i<n$. It is clear that we must add all the roots of $R_{n},n\in\mathbb{N}$ to $\zeta_{D,\gamma}$, and we cannot add anything else.
@@
Proposition 2.3. (Sufficient condition) Let $\lambda\in\mathfrak{n}^{*}$, $D=\textup{supp}(\lambda)$. Suppose that any root $\gamma\in D$ satisfy the following condition: if $\lambda([e_{\alpha},e_{\gamma}])\neq 0$ then $\alpha\in\zeta_{D,\gamma}$. Then $\lambda\in S$.
@@
Proof. Fix a root $\gamma\in D$. Consider the orbit of the tangent action $\mathfrak{n}.\lambda$. Let $\mu\in\mathfrak{n}.\lambda$, $\mu=x.\lambda$, $x=\sum\limits_{\alpha\in\Phi^{+}}x_{\alpha}e_{\alpha}$, while $\mu$ is given, and the values of $x_{\alpha}$ are not.
@@
We will show by induction on $|\zeta_{D,\gamma}|$ that if $\delta\in\zeta_{D,\gamma}$, then $x_{\delta}$ can be expressed as a linear combination of $\{\mu(e_{\beta})\}_{\beta\succ\gamma}$. Indeed, consider the step of adding $\delta$ to $\zeta_{D,\gamma}$. Suppose that we proved this for all roots that are already in $\zeta_{D,\gamma}$. There exists a root $\beta\succ\gamma$ such that $\delta$ is the unique root not from $\zeta_{D,\gamma}$ such that $\lambda([e_{\delta},e_{\beta}])\neq 0$. Consider $\mu(e_{\beta})=-\sum\limits_{\alpha\in\Phi^{+}}\lambda([e_{\alpha},e_{\beta}])x_{\alpha}$. Let us rewrite:
@@
We claim that the right hand side is a linear combination of $\{\mu(e_{\beta^{\prime}})\}_{\beta^{\prime}\succ\gamma}$. Indeed, if the coefficient at $x_{\alpha}$ in the right hand side is non-zero (i.e., $\lambda([e_{\alpha},e_{\beta}])\neq 0$), then $\alpha\in\zeta_{D,\gamma}$, hence, by the inductive assumption, $x_{\alpha}$ can be expressed as a linear combination of $\{\mu(e_{\beta^{\prime}})\}_{\beta^{\prime}\succ\gamma}$. Thus, since $\lambda([e_{\delta},e_{\beta}])\neq 0$, $x_{\delta}$ can also be expressed as a linear combination of $\{\mu(e_{\beta^{\prime}})\}_{\beta^{\prime}\succ\gamma}$.
@@
Note that if we take another form $\lambda^{\prime}$ with support $D$, then the conditions of Proposition Classification of coadjoint orbits for the maximal unipotent subgroup in the simple group of type $F_{4}$ will also be satisfied for it. So, for a given support $D$, either all forms with this support satisfy the conditions of Proposition Classification of coadjoint orbits for the maximal unipotent subgroup in the simple group of type $F_{4}$, or none. Therefore, it is sufficient to check the conditions for one form with a given support, for example, for the form with coordinates 1 at $e_{\alpha}$, where $\alpha$ lies in the support.

### v4

@@ Abstract
Let NN be the maximal unipotent subgroup of the simple algebraic group of type $\Phi$. It naturally acts on the space $\mathfrak{n}^{*}$ dual to the Lie algebra $\mathfrak{n}$ of NN, and this action is called the coadjoint action. Such orbits play a key role in A.A. Kirillov’s orbit method. In this work, we classify the orbits of this action in the case of $\Phi=F_{4}$ in terms of supports of canonical forms. This means that we will present a set $S$ of linear forms from $\mathfrak{n}^{*}$ such that for any coadjoint orbit there exists a unique form from $S$ belonging to that orbit. The set of canonical forms will be explicitly described in terms of supports. The support of a form $\lambda\in\mathfrak{n}^{*}$ is the set of positive roots $\alpha\in\Phi^{+}$ such that $\lambda(e_{\alpha})\neq 0$.
@@ 1. Introduction and the main result
Let $\mathfrak{g}$ be a simple finite-dimensional Lie algebra over a finite field $\mathbb{F}_{q}$ of sufficiently large characteristic $p$, $\mathfrak{b}$ be a Borel subalgebra of $\mathfrak{g}$, $\Phi$ be the root system of $\mathfrak{g}$, $\Phi^{+}$ be the set of positive roots corresponding to $\mathfrak{b}$, $\mathfrak{n}$ be the nilradical of $\mathfrak{b}$, $N=\exp(\mathfrak{n})$ be the corresponding unipotent algebraic group, and $\mathfrak{n}^{*}$ be the dual space of $\mathfrak{n}$. The group NN is also a maximal unipotent subgroup of the simple algebraic group $G$ of type $\Phi$, or equivalently a Sylow $p$-subgroup of $G$. The group NN acts on $\mathfrak{n}$ by the adjoint action; the dual action of NN on the space $\mathfrak{n}^{*}$ is called the coadjoint action; we will denote the result of this action by $g.\lambda$ for $g\in N$, $\lambda\in\mathfrak{n}^{*}$. According to the orbit method introduced by A.A. Kirillov in 1962, coadjoint orbits play the key role in the representation theory of the group NN. Almost all coadjoint orbits that have been studied well so far are so-called orbits associated with rook placements. They are discussed in detail in the works [An95], [AN06], [Ko12], [Ko13], [Pa08], [IPa09], [Ig09], [Ig11], [Ig12], [IS21] and [IS23].
Explicitly, the coadjoint action is given by:
@@
Let $\{e_{\alpha},\alpha\in\Phi^{+}\}$ be a basis of $\mathfrak{n}$ consisting of root vectors, and let $\{e_{\alpha}^{*},\alpha\in\Phi^{+}\}$ be the dual basis in $\mathfrak{n}^{*}$. The support of a linear form $\lambda\in\mathfrak{n}^{*}$ is defined as the subset of positive roots given by
@@
Our goal is to describe the coadjoint orbits for the case of $\Phi=F_{4}$. The description will be given in terms of supports of canonical forms. A list of supports will be provided with the following properties.
  • For any two linear forms $\lambda_{1}\neq\lambda_{2}$ whose supports belong to the given list (even if the supports coincide), the orbits $N.\lambda_{1}$ and $N.\lambda_{2}$ are different.
  • For any linear form $\lambda\in\mathfrak{n}^{*}$, there exists a form $\lambda^{\prime}$ with a support from the given list such that $N.\lambda=N.\lambda^{\prime}$.
In other words, each coadjoint orbit contains a unique linear form whose support belongs to the given list.
The set of simple roots of $F_{4}$ can be identified with the following subset of $\mathbb{R}^{4}$:
@@
Here $\{\varepsilon_{i}\}_{i=1}^{4}$ is the standard basis of $\mathbb{R}^{4}$ (endowed with the standard scalar product).
Recall that there exists a natural partial order on $\Phi^{+}$ defined as follows: $\alpha>\beta$ if $\alpha-\beta$ is a sum of positive roots. Let us fix a linear lexicographic order $\succ$ on the positive roots of $F_{4}$. We say that $\alpha=a_{1}\alpha_{1}+a_{2}\alpha_{2}+a_{3}\alpha_{3}+a_{4}\alpha_{4}\succ\beta=b_{1}\alpha_{1}+b_{2}\alpha_{2}+b_{3}\alpha_{3}+b_{4}\alpha_{4}$ if $a_{1}>b_{1}$ or $a_{1}=b_{1},a_{2}>b_{2}$ or $a_{1}=b_{1},a_{2}=b_{2},a_{3}>b_{3}$ or $a_{1}=b_{1},a_{2}=b_{2},a_{3}=b_{3},a_{4}>b_{4}$. Thus, the roots will be ordered as follows:
@@
This theorem generalizes the result of the work [IS23]. This article discusses the so-called orbits associated with rook placements. By definition, a subset $D$ of $\Phi^{+}$ is called a non-singular rook placement if $\alpha-\beta\notin\Phi^{+}$ for all distinct $\alpha,\beta\in D$. It follows that $(\alpha,\beta)\leq 0$ for all distinct $\alpha,\beta\in D$. Given a rook placement $D$ and a map $\xi\colon D\to\mathbb{C}^{\times}$, we put
@@
We say that the coadjoint orbit $\Omega_{D,\xi}$ of the linear form $f_{D,\xi}$ is associated with the rook placement $D$ and the map $\xi$.Thus, $f_{D,\xi}$ is a linear form with support $D$.
In [IS23] it is proved that if $\Phi=F_{4}$, $D$ is an orthogonal non-singular rook placement and $\xi_{1}$, $\xi_{2}$ are distinct maps from $D$ to $\mathbb{C}^{\times}$ then $\Omega_{D,\xi_{1}}\neq\Omega_{D,\xi_{2}}$. Moreover, if $D$ is a non-singular rook placement, then $\lambda=f_{D,\xi}$ belongs to $S$. Indeed, if $\gamma\in D$, then $\gamma+\beta\notin D$ for all $\beta\in\Phi^{+}$ and $\lambda([e_{\gamma},e_{\beta}])=0$ for all $\beta\in\Phi^{+}$. Hence, in the matrix $A_{\lambda,\gamma}$ the row corresponding to $\gamma$ is zero, and condition (1) is true for $\gamma\in D$. Therefore, distinct linear forms with the same support $D$ lie in different coadjoint orbits.
A result similar to Theorem Classification of coadjoint orbits for the maximal unipotent subgroup in the simple group of type $F_{4}$ holds for the root system $\Phi=G_{2}$ and field $\mathbb{C}$. In the same work [IS23], the equations defining the basic subvarieties $\mathcal{O}_{D,\xi}$ are given. By definition, the basic subvariety $\mathcal{O}_{D,\xi}$ corresponding to a rook placement $D$ and a map $\xi\colon D\to\mathbb{C}^{\times}$ is
@@
where $\xi_{\alpha}$ is the restriction of $\xi$ to $\{\alpha\}$. That work also proves AndrГ©’s stratification for $G_{2}$:
@@
For almost every rook placement $D$, we have $\mathcal{O}_{D,\xi}=\Omega_{D,\xi}$. The only exception is the case $D=\{\alpha+\beta,3\alpha+\beta\}$. In this case $\mathcal{O}_{D,\xi}$ is union of orbits $\Omega_{D,\xi}$ and $\Omega_{\{\beta,\alpha+\beta,3\alpha+\beta\},\xi^{\prime}}$ where $\xi^{\prime}\colon{\beta,\alpha+\beta,3\alpha+\beta}\to\mathbb{C}^{\times}$ is such that its restriction to $D$ coincides with $\xi$. Let us write down the equations defining the orbits for $D={\alpha+\beta,3\alpha+\beta}$. A linear form $\lambda=\sum_{\gamma\in\Phi^{+}}\lambda_{\gamma}e_{\gamma}^{*}$ lies in $\Omega_{D,\xi}$ if and only if it satisfies the following system:
@@
The orbits of these forms correspond precisely to all the orbits whose equations are provided in [IS23] and in the present work.
@@
Acknowledgments. I thank M.V. Ignatev and A.N. Panov for helpful comments and support during the writing of this paper.
@@ 2. Description of the set of canonical forms
In this section, we will obtain an explicit description of the set $S$ to prove Theorem Classification of coadjoint orbits for the maximal unipotent subgroup in the simple group of type $F_{4}$ and to demonstrate its use. At first glance, it might seem that this set does not admit a nice description. In fact, this is not the case. The set $S$ can be described very neatly in terms of supports. A list of 880 supports is presented in the Appendix.
@@
and of the forms with one of the three supports listed below, provided their coordinates satisfy the equation
@@
Here $\lambda=\sum_{\alpha\in\Phi^{+}}\lambda_{\alpha}e_{\alpha}^{*}$, and $\lambda_{\alpha}$ are the coordinates with respect to the dual basis.
@@
In what follows, we will use the following table. It is a $24\times 24$ table whose rows and columns are indexed by the positive roots. In the cell at the intersection of row $\alpha$ and column $\beta$, we write $\beta-\alpha$ if $\beta-\alpha$ can be expressed as a sum of positive roots; otherwise, the cell is left blank. The main diagonal contains zeros. It is easy to see that all cells below the diagonal are empty. Below we show the upper-left corner of this table, corresponding to those positive roots whose simple root expansion does not contain $\alpha_{1}$. A cell at position $(\alpha,\beta)$ is highlighted in orange if the difference $\beta-\alpha$ is itself a positive root. Almost all examples will involve specifically this part of the table. Four columns are highlighted in blue, as they will appear most frequently in the examples.
@@
To describe the set $S$, we wrote a program, which is included in the Appendix. The program outputs supports that satisfy a sufficient condition for being included in the list, and supports that satisfy a necessary condition; these conditions will be described below.
First, let us reformulate condition (1). The algebra $\mathfrak{n}$ acts on $\mathfrak{n}^{*}$ as follows: for $x,y\in\mathfrak{n}$ and $\lambda\in\mathfrak{n}^{*}$, define $(x.\lambda)(y):=-\lambda([x,y])$. This action is called tangent.
Let us expand $x\in\mathfrak{n}$ in the chosen basis:
@@
We can view these coordinates as linear combinations of variables $x_{\alpha},\alpha\in\Phi^{+}$, with constant coefficients $\lambda([e_{\alpha},e_{\beta}])$. By definition, $A_{\lambda,\gamma}$ is the matrix of linear combinations $\{(x.\lambda)(e_{\alpha})\}_{\alpha\succeq\gamma}$, and $B_{\lambda,\gamma}$ is the matrix of linear combinations $\{(x.\lambda)(e_{\alpha})\}_{\alpha\succ\gamma}$. Clearly, $\mathrm{rk}\,A_{\lambda,\gamma}=\mathrm{rk}\,B_{\lambda,\gamma}$ if and only if the row numbered $\gamma$ of the matrix $A_{\lambda,\gamma}$ is a linear combination of all its other rows (i.e., the rows of the matrix $B_{\lambda,\gamma}$). In other words, $(x.\lambda)(e_{\gamma})$ can be expressed as a linear combination of $\{(x.\lambda)(e_{\alpha})\}_{\alpha\succ\gamma}$.
Example 2.2. Let $D=\textup{supp}(\lambda)=\{\alpha_{4},\alpha_{3}+\alpha_{4},\alpha_{2},\alpha_{2}+2\alpha_{3}\}$. We will show that such a $\lambda$ lies in $S$. To visualize what the orbit of the tangent action looks like, let us refer to the table. The highlighted blue columns correspond to the roots in $D$. The linear combination $(x.\lambda)(e_{\beta})$ contains only those $x_{\alpha}$ for which $\lambda([e_{\alpha},e_{\beta}])\neq 0$, which is equivalent to $\alpha+\beta\in D$. It can be seen from the construction of the table that $(x.\lambda)(e_{\beta})$ contains only $x_{\alpha}$ such that the orange root $\alpha$ is in the row corresponding to $\beta$ in the highlighted column. For example, the combination $(x.\lambda)(e_{\alpha_{3}})$ contains only $x_{\alpha_{4}}$ and $x_{\alpha_{2}+\alpha_{3}}$.
@@
It is clear that we did not need to know the coefficients for $x_{\alpha_{3}}$ in either $(x.\lambda)(e_{\alpha_{2}+\alpha_{3}})$ or $(x.\lambda)(e_{\alpha_{4}})$. This is simply because $x_{\alpha_{3}}$ as a variable can be expressed in terms of $(x.\lambda)(e_{\alpha_{2}+\alpha_{3}})$, and everywhere above (i.e., in $(x.\lambda)(e_{\alpha_{2}})$, $(x.\lambda)(e_{\alpha_{3}+\alpha_{4}})$, $(x.\lambda)(e_{\alpha_{3}})$, $(x.\lambda)(e_{\alpha_{4}})$) $x_{\alpha_{3}}$ can be expressed as $(x.\lambda)(e_{\alpha_{2}+\alpha_{3}})$ multiplied by some coefficient. The Sufficient condition below is based on this idea.
@@
Consider the tangent action $\mathfrak{n}\curvearrowright\mathfrak{n}^{*}$. Let $\gamma\in\Phi^{+}$, $\lambda\in\mathfrak{n}^{*}$, $D=\textup{supp}(\lambda)$. Starting from the support $D$ and the root $\gamma$, we will construct a set $\zeta_{D,\gamma}$ of positive roots. We will add roots to $\zeta_{D,\gamma}$ one by one. We add a root $\delta$ to $\zeta_{D,\gamma}$ if there exists a root $\beta\succ\gamma$ such that $\lambda([e_{\delta},e_{\beta}])\neq 0$ and if $\lambda([e_{\alpha},e_{\beta}])\neq 0$ for some other $\alpha\in\Phi^{+}$, then $\alpha\in\zeta_{D,\gamma}$. Using this procedure, we will add to $\zeta_{D,\gamma}$ the maximum possible number of roots. Note that $\zeta_{D,\gamma}$ does not depend on the specific values of the coordinates of $\lambda$ in the dual basis, but only on its support $D$. The set $\zeta_{D,\gamma}$ also does not depend on the order in which we add the roots. Indeed, define sets $R_{n}$ for $n\in\mathbb{N}$. A root $\delta$ lies in $R_{n}$ if there exists a root $\beta\succ\gamma$ such that $\lambda([e_{\delta},e_{\beta}])\neq 0$ and if $\lambda([e_{\alpha},e_{\beta}])\neq 0$ for some other $\alpha\in\Phi^{+}$, then $\alpha\in R_{i}$ for some $i<n$. Clearly, we must add all roots from $\bigcup_{n\in\mathbb{N}}R_{n}$ to $\zeta_{D,\gamma}$, and we cannot add any other roots.
@@
Proposition 2.3. (Sufficient condition) Let $\lambda\in\mathfrak{n}^{*}$, $D=\operatorname{supp}(\lambda)$. Suppose that for every root $\gamma\in D$, the following condition holds: if $\lambda([e_{\alpha},e_{\gamma}])\neq 0$, then $\alpha\in\zeta_{D,\gamma}$. Then $\lambda\in S$.
@@
Proof. Fix a root $\gamma\in D$. Consider the orbit of the infinitesimal action $\mathfrak{n}.\lambda$. Let $\mu\in\mathfrak{n}.\lambda$, $\mu=x.\lambda$, where $x=\sum_{\alpha\in\Phi^{+}}x_{\alpha}e_{\alpha}$. Here $\mu$ is given, while the values $x_{\alpha}$ are unknown.
@@
We will show by induction on $|\zeta_{D,\gamma}|$ that if $\delta\in\zeta_{D,\gamma}$, then $x_{\delta}$ can be expressed as a linear combination of $\{\mu(e_{\beta})\}_{\beta\succ\gamma}$. Consider the step of adding $\delta$ to $\zeta_{D,\gamma}$. Suppose we have already proved this for all roots that are already in $\zeta_{D,\gamma}$. There exists a root $\beta\succ\gamma$ such that $\delta$ is the unique root not in $\zeta_{D,\gamma}$ satisfying $\lambda([e_{\delta},e_{\beta}])\neq 0$. Consider $\mu(e_{\beta})=-\sum_{\alpha\in\Phi^{+}}\lambda([e_{\alpha},e_{\beta}])x_{\alpha}$. This can be rewritten as:
@@
We claim that the right-hand side is a linear combination of $\{\mu(e_{\beta^{\prime}})\}_{\beta^{\prime}\succ\gamma}$. Indeed, if the coefficient of $x\alpha$ on the right-hand side is non-zero (i.e., $\lambda([e_{\alpha},e_{\beta}])\neq 0$), then $\alpha\in\zeta_{D,\gamma}$. By the inductive hypothesis, $x_{\alpha}$ can be expressed as a linear combination of $\{\mu(e_{\beta^{\prime}})\}_{\beta^{\prime}\succ\gamma}$. Since $\lambda([e_{\delta},e_{\beta}])\neq 0$, it follows that $x_{\delta}$ can also be expressed as a linear combination of $\{\mu(e_{\beta^{\prime}})\}_{\beta^{\prime}\succ\gamma}$.
@@
Note that if we take another form $\lambda^{\prime}$ with support $D$, then the conditions of Proposition Classification of coadjoint orbits for the maximal unipotent subgroup in the simple group of type $F_{4}$ will also be satisfied for it. Therefore, for a given support $D$, either all forms with this support satisfy the conditions of Proposition Classification of coadjoint orbits for the maximal unipotent subgroup in the simple group of type $F_{4}$, or none do. Hence, it suffices to check the conditions for a single form with a given support, for example, for the form with all coordinates equal to 11 at the basis vectors $e_{\alpha}^{*}$ with $\alpha\in D$.